# Manipulating 3D Molecules in a Fixed-Dimensional SE(3)-Equivariant Latent Space

### v1

@@ Abstract
Medicinal chemists often optimize drugs considering their 3D structures and designing structurally distinct molecules that retain key features, such as shapes, pharmacophores, or chemical properties. Previous deep learning approaches address this through supervised tasks like molecule inpainting or property-guided optimization. In this work, we propose a flexible zero-shot molecule manipulation method by navigating in a shared latent space of 3D molecules. We introduce a Variational AutoEncoder (VAE) for 3D molecules, named MolFLAE, which learns a fixed-dimensional, SE(3)-equivariant latent space independent of atom counts. MolFLAE encodes 3D molecules using an SE(3)-equivariant neural network into fixed number of latent nodes, distinguished by learned embeddings. The latent space is regularized, and molecular structures are reconstructed via a Bayesian Flow Network (BFN) conditioned on the encoder’s latent output. MolFLAE achieves competitive performance on standard unconditional 3D molecule generation benchmarks. Moreover, the latent space of MolFLAE enables zero-shot molecule manipulation, including atom number editing, structure reconstruction, and coordinated latent interpolation for both structure and properties. We further demonstrate our approach on a drug optimization task for the human glucocorticoid receptor, generating molecules with improved hydrophilicity while preserving key interactions, under computational evaluations. These results highlight the flexibility, robustness, and real-world utility of our method, opening new avenues for molecule editing and optimization.
@@ 1 Introduction
To address these challenges, we propose MolFLAE (Molecule Fixed Length AutoEncoder), a Variational AutoEncoder (VAE) for 3D molecules that learns a fixed-dimensional, SE(3)-equivariant latent space, independent of atom counts. Our encoder employs an SE(3)-equivariant neural network that updates a fixed number of virtual nodes initialized with learnable embeddings, transforming them into fixed-length latent codes for 3D molecules. The latent space is regularized under the standard VAE framework, and a Bayesian Flow Network (BFN) serves as the decoder, reconstructing full molecular structures conditioned on latent codes.
@@
  • We propose a VAE model that learns a fixed-dimensional, SE(3)-equivariant latent space for 3D molecular structures;
@@ 2 Related Works
Unconditional Generation for 3D Molecules  Unconditional 3D molecule generation has achieved rapid advancements, driven by progress in deep generative models. Early works explored auto-regressive models to construct molecules atom-by-atom (luo2022an, ; DBLP:conf/nips/GebauerGS19, ). More recently, inspired by diffusion-like models (ho2020denoising, ; song2021scorebased, ; DBLP:conf/iclr/LipmanCBNL23, ; DBLP:conf/iclr/LiuG023, ; graves2023bayesian, ), models like EDM (DBLP:conf/icml/HoogeboomSVW22, ), EquiFM (song2023equivariant, ), and GeoBFN (song2024unified, ), have significantly improved generation quality. VAEs (DBLP:journals/corr/KingmaW13, ) offer an alternative by decoding latent embeddings into molecules, but modeling 3D molecules is challenging due to variable atom counts and the need for equivariance of coordinates to rotations and translations. As a result, existing 3D molecular VAEs (DBLP:conf/icml/XuPDEL23, ; kong2024fullatom, ) often encode atoms into separate latent vectors, yielding variable-length latent representations.
@@
Generate 3D Molecules from Fixed-Dimensional Latent Spaces Fixed-length autoencoders for molecules have been explored through voxel-based models, which discretize 3D space into uniform grids and apply 3D CNNs (masuda2020generating, ; pinheiro2023d, ) or neural fields (kirchmeyer2024scorebased, ). While straightforward, these methods are not SE(3)-equivariant and struggle to disentangle semantic features from orientations, making molecule manipulation in the latent space not flexible. Local-frame-based models (Winter2021AutoEncodingMC, ) represent conformations with SE(3) invariant features, inculding distances, bond angles, and dihedrals, making outputs orientation-agnostic. UAE-3D (luo2025unifiedlatentspace3d, ) has proposed a 3D fixed-length latent space by discarding the inductive bias of geometric equivariance, yet their latent space suffers from similar feature entanglement problem as voxel-based models. A recent work (li2024d, ) also constructs fixed-length autoencoder, but applies global pooling over atom features, discarding spatial information that is critical for reconstruction and interaction modeling.
@@ 3 Methodology
We encode a variable-size 3D molecule by concatenating it with a set of learnable virtual nodes and updating them together via an SE(3)-equivariant neural network to obtain fixed-length embeddings. The regularization loss $\mathcal{L}_{\text{reg}}$ comes from an Multi-Layer Perceptron (MLP) predicting the means and variances to form a variational posterior, from which we sample latent codes that condition a Bayesian Flow Network (BFN) decoder for reconstruction producing the reconstruction loss $\mathcal{L}_{\text{recon}}$. The MolFLAE is trained end-to-end by minimizing
@@ 3.3 MolFLAE Encoder
We employ an SE(3)-equivariant neural network $\bm{\phi}_{\theta}$ to jointly encode the original molecular point cloud and the appended virtual nodes $(\mathcal{M},\mathcal{Z})$. After rounds of update, we discard the embeddings of $\mathcal{M}$ and retain only those of $\mathcal{Z}$ as our fixed-length latent representation. In explicit,
(_$(\ \_\ ,[\mathbf{z}_{x},\mathbf{z}_{h}])=\bm{\phi}_{\theta}\left([\mathbf{x}_{M},\mathbf{v}_{M}],[\mathbf{x}_{Z},\mathbf{v}_{Z}]\right).$ (3)
The latent code is denoted as $[\mathbf{z}_{x},\mathbf{z}_{h}]\in\mathbb{R}^{N_{Z}\times(3+D_{f})}$, where $\mathbf{z}_{x}$ and $\mathbf{z}_{h}$ represent the spatial and feature components, $D_{f}$ is the embedding dimension of features. Note that we only keep the fixed-length part of the $\phi_{\theta}$ output. So we obtain a fixed-length encoding of the molecules. The network structure and SE(3)-equivariance discussion can be found in the Appendix A.
To regularize the latent space, we adopt a VAE formulation. While the initial output $[\mathbf{z}_{x},\mathbf{z}_{h}]$ is deterministic, this can lead to irregular latent geometry and poor interpolatability. To address this, we predict a coordinate-wise Gaussian distribution for each latent dimension:
$\bm{\mu}_{x}=\mathbf{z}_{x},\quad[\bm{\sigma}^{2}_{x},\;\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]=\mathtt{Linear}([\mathbf{z}_{x},\mathbf{z}_{h}]).$ (4)
@@ 3.4 MolFLAE Decoder
Then we can derive the two loss $\mathcal{L}_{x}^{n}+\mathcal{L}_{v}^{n}$ respectively (qu2024molcraft, ). The total forward pass can be found in the Alogorithm. 2. The inference process is parallel with the general BFN inference but taking two data modelities into consideration. See Appendix C.4
@@ 4 Experiments
To assess the capability of MolFLAE generate stable, diverse molecules, we first focus on 3D molecule generation task following the setting of prior works (DBLP:conf/icml/HoogeboomSVW22, ; song2023equivariant, ; song2024unified, ). We conduct 10,000 random samplings in the latent space, then decode them into molecules using MolFLAE decoder, subsequently evaluating qualities of these molecules. We sample the atom number from the prior of the training set as previous works like (DBLP:conf/icml/HoogeboomSVW22, ). Table 1 illustrates the benchmark results of unconditional generation with MolFLAE.
@@
The latent space of MolFLAE consists of two parts, the SE(3)-equivariant component $\mathbf{z}_{x}$ and the SE(3)-invariant component $\mathbf{z}_{h}$. Ideally, spatial and semantic features of molecules disentangle spontaneously, with $\mathbf{z}_{x}$ encoding the shape and orientation and $\mathbf{z}_{h}$ encodes substructures of molecules.
@@
In experiments, we have observed a partial disentanglement of the MolFLAE latent space. As shown in Fig. 3, following the disentanglement hypothesis, substructures information of $M_{1}$ should be able to be extracted by isolating $\mathbf{z}_{x}^{1}$. As $\mathbf{z}_{x}^{1}$ is not sufficient to reconstruct $M_{1}$, we view it as a deconstructed molecule. Then, we reconstruct these substructures into the shape and orientation of $M_{0}$, by decoding the hybrid latent code $\left(\mathbf{z}_{h}^{1},\mathbf{z}_{x}^{0}\right)$. The resulted molecule shares a similar shape and orientation with $M_{0}$, indicated by the dash lines in Fig 3. Moreover, it also shares similar substructures as $M_{1}$ like amide, chlorobenzyl, sulfamide, indicated by rectangles of corresponding colors.
@@
Metrics MACCS Sim $\downarrow$ Shape Sim $\uparrow$ MACCS Sim $\uparrow$ Shape Sim $\downarrow$
MolFLAE 0.421 0.394 0.580 0.174
@@
To further quantify the quality of interpolation, We evaluate the trend of properties of the intermediate molecules during the transformation, considering both structural and physical properties (detailed descriptions are provided in Appendix D). We compute the Pearson correlation coefficient $r$ between each property value and its corresponding interpolation index (adjusted with the sign of property difference between the source and target molecule). We also report the associated $p$-values from the Pearson significance test to assess the statistical evidence for linear trends. A property is considered to exhibit significant linear variation along the interpolation trajectory if the null hypothesis of zero correlation is rejected at the 5% significance level, i.e., if $-\log p>-\log(0.05)\approx 1.3$. Detailed results are documented in Table 5 and Table 4.
@@ 5 Conclusion and Future Works
In this work, we present MolFLAE, a flexible VAE framework for manipulating 3D molecules within a fixed-dimensional, SE(3)-equivariant latent space. Our method demonstrates strong performance across multiple tasks, including unconditional generation, analog design, substructure reconstruction, and latent interpolation. We further validate the real-world utility of MolFLAE through a case study on generating drug-like molecules targeting the hGR, balancing potency and solubility.
@@
In summary, our results highlight the versatility of MolFLAE and its promise as a general-purpose framework for 3D molecule generation and editing. This work opens new directions for exploring the broader applications of fixed-dimensional, SE(3)-equivariant latent spaces in molecular modeling.
@@ Appendix A Neural Network Details
We adopt the network structure of MolCRAFT [36] as our SE(3)-Equivariant graph neural network (GNN) backbone. Originally designed to model interactions between ligand and protein pocket atoms, the network distinguishes between update nodes $\mathbf{u}$ (whose coordinates are updated) and condition nodes $\mathbf{c}$ (which provide contextual information and the coordinates are not updated).
@@ A.1 Neural Network Architecture
To be convenient, we concatenate the spatial part and feature part of the update nodes $[\mathbf{x}_{\text{update}},\mathbf{v}_{\text{update}}]$ and condition nodes $[\mathbf{x}_{\text{condition}},\mathbf{v}_{\text{condition}}]$ with writing $\mathbf{x}^{\ell}=[\mathbf{x}_{\text{update}}^{\ell},\mathbf{x}_{\text{condition}}^{\ell}]$ and $\mathbf{h}^{\ell}=[\mathbf{v}_{\text{update}}^{\ell},\mathbf{v}_{\text{condition}}^{\ell}]$, where the superscript represents the $\ell$-th layer of $\bm{\phi}$, $0\leq\ell\leq L$. The Initial hidden embedding $\mathbf{h}^{0}$ is obtained by an MLP embedding layer that encodes the atom feature $[\mathbf{h}]$. No embedding layer for the atom spatial coordinates. The construction of $\bm{\phi}_{\theta}$ is alternately updating the atom feature embeddings $\mathbf{h}$ and coordinates $\mathbf{x}$ as
@@
[…]
@@
A.2 SE(3)-Equivariance Discussion
In molecular modeling, it is essential that the learned distribution over update nodes be invariant to translations and rotations of the condition nodes. This SE(3)-invariance reflects a fundamental inductive bias in molecular systems [15, 47].
The Special Euclidean group in $\mathbb{R}^{3}$, denoted SE(3), consists of rigid-body transformations of the form $T(x)=Rx+t$, where $R\in\mathbb{R}^{3\times 3}$ is a rotation matrix and $t\in\mathbb{R}^{3}$ is a translation vector.
If we pre-align the condition nodes by centering them at their center of mass (i.e., eliminating the translational degree of freedom), then the resulting likelihood becomes SE(3)-invariant under the following condition:

### v2

# Manipulating 3D Molecules in a Fixed-Dimensional E(3)-Equivariant Latent Space

@@ Abstract
Medicinal chemists often optimize drugs considering their 3D structures and designing structurally distinct molecules that retain key features, such as shapes, pharmacophores, or chemical properties. Previous deep learning approaches address this through supervised tasks like molecule inpainting or property-guided optimization. In this work, we propose a flexible zero-shot molecule manipulation method by navigating in a shared latent space of 3D molecules. We introduce a Variational AutoEncoder (VAE) for 3D molecules, named MolFLAE, which learns a fixed-dimensional, E(3)-equivariant latent space independent of atom counts. MolFLAE encodes 3D molecules using an E(3)-equivariant neural network into fixed number of latent nodes, distinguished by learned embeddings. The latent space is regularized, and molecular structures are reconstructed via a Bayesian Flow Network (BFN) conditioned on the encoder’s latent output. MolFLAE achieves competitive performance on standard unconditional 3D molecule generation benchmarks. Moreover, the latent space of MolFLAE enables zero-shot molecule manipulation, including atom number editing, structure reconstruction, and coordinated latent interpolation for both structure and properties. We further demonstrate our approach on a drug optimization task for the human glucocorticoid receptor, generating molecules with improved hydrophilicity while preserving key interactions, under computational evaluations. These results highlight the flexibility, robustness, and real-world utility of our method, opening new avenues for molecule editing and optimization.  [^1] The code is available at https://github.com/MuZhao2333/MolFLAE
@@ 1 Introduction
To address these challenges, we propose MolFLAE (Molecule Fixed Length AutoEncoder), a Variational AutoEncoder (VAE) for 3D molecules that learns a fixed-dimensional, E(3)-equivariant latent space, independent of atom counts. Our encoder employs an E(3)-equivariant neural network that updates a fixed number of virtual nodes initialized with learnable embeddings, transforming them into fixed-length latent codes for 3D molecules. The latent space is regularized under the standard VAE framework, and a Bayesian Flow Network (BFN) serves as the decoder, reconstructing full molecular structures conditioned on latent codes.
@@
  • We propose a VAE model that learns a fixed-dimensional, E(3)-equivariant latent space for 3D molecular structures;
@@ 2 Related Works
Unconditional Generation for 3D Molecules  Unconditional 3D molecule generation has achieved rapid advancements, driven by progress in deep generative models. Early works explored auto-regressive models to construct molecules atom-by-atom (luo2022an, ; DBLP:conf/nips/GebauerGS19, ). More recently, inspired by diffusion-like models (ho2020denoising, ; song2021scorebased, ; DBLP:conf/iclr/LipmanCBNL23, ; DBLP:conf/iclr/LiuG023, ; graves2023bayesian, ), models like EDM (DBLP:conf/icml/HoogeboomSVW22, ), EquiFM (song2023equivariant, ), and GeoBFN (song2024unified, ), have significantly improved generation quality. VAEs (DBLP:journals/corr/KingmaW13, ) offer an alternative by decoding latent embeddings into molecules, but modeling 3D molecules is challenging due to variable atom counts and the need for equivariance of coordinates to rotations and translations. Existing 3D molecular VAEs such as PepGLAD (kong2024fullatom, ) and UniMoMo (kong2025unimomo, ) have made progress by encoding molecular blocks (e.g., amino acid residues) into independent latent nodes. This design ensures E(3)-equivariance but has two key limitations: (1) the number of latent nodes depends on the molecular composition, complicating cross-sample comparisons and interpolation tasks; and (2) spatial relationships between molecular blocks are tightly preserved, restricting the flexibility of the latent space for generative modeling, especially when latent codes are repurposed for tasks like autoregressive modeling by finetuning LLMs.
@@
Generate 3D Molecules from Fixed-Dimensional Latent Spaces Fixed-length autoencoders for molecules have been explored through voxel-based models, which discretize 3D space into uniform grids and apply 3D CNNs (masuda2020generating, ; pinheiro2023d, ) or neural fields (kirchmeyer2024scorebased, ). While straightforward, these methods are not E(3)-equivariant and struggle to disentangle semantic features from orientations, making molecule manipulation in the latent space not flexible. Local-frame-based models (Winter2021AutoEncodingMC, ) represent conformations with SE(3) invariant features, including distances, bond angles, and dihedrals, making outputs orientation-agnostic. UAE-3D (luo2025unifiedlatentspace3d, ) has proposed a 3D fixed-length latent space by discarding the inductive bias of geometric equivariance, yet their latent space suffers from similar feature entanglement problem as voxel-based models. A recent work (li2024d, ) also constructs fixed-length autoencoder, but applies global pooling over atom features, discarding spatial information that is critical for reconstruction and interaction modeling.
@@ 3 Methodology
We encode a variable-size 3D molecule by concatenating it with a set of learnable virtual nodes and updating them together via an E(3)-equivariant neural network to obtain fixed-length embeddings. The regularization loss $\mathcal{L}_{\text{reg}}$ comes from an Multi-Layer Perceptron (MLP) predicting the means and variances to form a variational posterior, from which we sample latent codes that condition a Bayesian Flow Network (BFN) decoder for reconstruction producing the reconstruction loss $\mathcal{L}_{\text{recon}}$. The MolFLAE is trained end-to-end by minimizing
@@ 3.3 MolFLAE Encoder
We employ an E(3)-equivariant neural network $\bm{\phi}_{\theta}$ to jointly encode the original molecular point cloud and the appended virtual nodes $(\mathcal{M},\mathcal{Z})$. After rounds of update, we discard the embeddings of $\mathcal{M}$ and retain only those of $\mathcal{Z}$ as our fixed-length latent representation. In explicit,
$(\ \_\ ,[\mathbf{z}_{x},\mathbf{z}_{h}])=\bm{\phi}_{\theta}\left([\mathbf{x}_{M},\mathbf{v}_{M}],[\mathbf{x}_{Z},\mathbf{v}_{Z}]\right).$ (3)
The latent code is denoted as $\mathbf{z}_{x},\mathbf{z}_{h}]\in\mathbb{R}^{N_{Z}\times(3+D_{f})}$, where $\mathbf{z}_{x}$ and $\mathbf{z}_{h}$ represent the spatial and feature components, $D_{f}$ is the embedding dimension of features. Note that we only keep the fixed-length part of the $\phi_{\theta}$ output. So we obtain a fixed-length encoding of the molecules. The network structure and E(3)-equivariance discussion can be found in the Appendix A.
To regularize the latent space, we adopt a VAE formulation. While the initial output $\mathbf{z}_{x},\mathbf{z}_{h}]$ is deterministic, this can lead to irregular latent geometry and poor interpolatability. To address this, we predict a coordinate-wise Gaussian distribution for each latent dimension:
$\bm{\mu}_{x}=\mathbf{z}_{x},\quad[\bm{\sigma}^{2}_{x},\;\bm{\mu}_{h},\bm{\sigma}^{2}_{h}]=\mathtt{Linear}(\mathbf{z}_{h}).$ (4)
@@ 3.4 MolFLAE Decoder
Then we can derive the two loss $\mathcal{L}_{x}^{n}+\mathcal{L}_{v}^{n}$ respectively (qu2024molcraft, ). The total forward pass can be found in the Algorithm. 2. The inference process is parallel with the general BFN inference but taking two data modalities into consideration. See Appendix C.4
@@ 4 Experiments
To assess the capability of MolFLAE generate stable, diverse molecules, we first focus on 3D molecule generation task following the setting of prior works (DBLP:conf/icml/HoogeboomSVW22, ; song2023equivariant, ; song2024unified, ). We conduct 10,000 random samplings in the latent space, then decode them into molecules using MolFLAE decoder, subsequently evaluating qualities of these molecules. We sample the atom number from the prior of the training set as previous works like (DBLP:conf/icml/HoogeboomSVW22, ). Table 1 illustrates the benchmark results of unconditional generation with MolFLAE. We also provide results on drug-likeness metrics on GEOM-Drugs in Appendix E, confirming MolFLAE’s outstanding performance in generating structurally reasonable and drug-like molecules compared to previous methods.
@@
The latent space of MolFLAE consists of two parts, the E(3)-equivariant component $\mathbf{z}_{x}$ and the E(3)-invariant component $\mathbf{z}_{h}$. Ideally, spatial and semantic features of molecules disentangle spontaneously, with $\mathbf{z}_{x}$ encoding the shape and orientation and $\mathbf{z}_{h}$ encodes substructures of molecules.
@@
In experiments, we have observed a partial disentanglement of the MolFLAE latent space. As shown in Fig. 3, following the disentanglement hypothesis, substructures information of $M_{1}$ should be able to be extracted by isolating $\mathbf{z}_{h}^{1}$. As $\mathbf{z}_{h}^{1}$ is not sufficient to reconstruct $M_{1}$, we view it as a deconstructed molecule. Then, we reconstruct these substructures into the shape and orientation of $M_{0}$, by decoding the hybrid latent code $\left(\mathbf{z}_{h}^{1},\mathbf{z}_{x}^{0}\right)$. The resulted molecule shares a similar shape and orientation with $M_{0}$, indicated by the dash lines in Fig 3. Moreover, it also shares similar substructures as $M_{1}$ like amide, chlorobenzyl, sulfamide, indicated by rectangles of corresponding colors.
@@
Metrics MACCS Sim $\downarrow$ Shape Sim $\uparrow$ Valid(%)($\uparrow$) Atom Sta(%)($\uparrow$) MACCS Sim $\uparrow$ Shape Sim $\downarrow$ Valid(%)($\uparrow$) Atom Sta(%)($\uparrow$)
MolFLAE 0.421 0.394 100.0 85.20 0.580 0.174 100.0 84.62
@@
To further quantify the quality of interpolation, We evaluate the trend of properties of the intermediate molecules during the transformation, considering both structural and physical properties (detailed descriptions are provided in Appendix D). We compute the Pearson correlation coefficient $r$ between each property value and its corresponding interpolation index (adjusted with the sign of property difference between the source and target molecule). We also report the associated $p$-values from the Pearson significance test to assess the statistical evidence for linear trends. A property is considered to exhibit significant linear variation along the interpolation trajectory if the null hypothesis of zero correlation is rejected at the 5% significance level, i.e., if $\log p>-\log(0.05)\approx 1.3$. Detailed results are documented in Table 5 and Table 4. We also report the step-wise molecular validity and atom stability during interpolation in Table 6, which indicates that most intermediate molecules are valid and stable.
@@
Table 6: Step-wise Validity and Atom Stability during interpolation.
Interpo Num Metrics Step
1 2 3 4 5 6 7 8 9 10 11 12
8 Valid(%) 99.90 99.90 99.90 100.0 99.90 100.0 100.0 100.0
Atom Sta(%) 82.23 82.95 84.17 84.57 84.21 83.74 83.10 82.63
10 Valid(%) 100.0 99.90 99.90 99.80 99.80 100.0 100.0 100.0 99.80 99.90
Atom Sta(%) 82.23 82.75 83.53 84.07 84.36 84.35 84.16 84.17 82.78 82.67
12 Valid(%) 99.90 100.0 99.90 100.0 100.0 100.0 100.0 100.0 99.90 99.90 100.0 99.80
Atom Sta(%) 82.39 82.50 83.76 83.67 83.92 84.65 84.20 84.50 84.27 83.00 82.84 82.57
@@ 5 Conclusion and Future Works
In this work, we present MolFLAE, a flexible VAE framework for manipulating 3D molecules within a fixed-dimensional, E(3)-equivariant latent space. Our method demonstrates strong performance across multiple tasks, including unconditional generation, analog design, substructure reconstruction, and latent interpolation. We further validate the real-world utility of MolFLAE through a case study on generating drug-like molecules targeting the hGR, balancing potency and solubility.
@@
In summary, our results highlight the versatility of MolFLAE and its promise as a general-purpose framework for 3D molecule generation and editing. This work opens new directions for exploring the broader applications of fixed-dimensional, E(3)-equivariant latent spaces in molecular modeling.
@@ Appendix A Neural Network Details
We adopt the network structure of MolCRAFT [37] as our E(3)-Equivariant graph neural network (GNN) backbone. Originally designed to model interactions between ligand and protein pocket atoms, the network distinguishes between update nodes $\mathbf{u}$ (whose coordinates are updated) and condition nodes $\mathbf{c}$ (which provide contextual information and the coordinates are not updated).
@@ A.1 Neural Network Architecture
To be convenient, we concatenate the spatial part and feature part of the update nodes $\mathbf{x}_{\text{update}},\mathbf{v}_{\text{update}}]$ and condition nodes $\mathbf{x}_{\text{condition}},\mathbf{v}_{\text{condition}}]$ with writing $\mathbf{x}^{\ell}=[\mathbf{x}_{\text{update}}^{\ell},\mathbf{x}_{\text{condition}}^{\ell}]$ and $\mathbf{h}^{\ell}=[\mathbf{v}_{\text{update}}^{\ell},\mathbf{v}_{\text{condition}}^{\ell}]$, where the superscript represents the $\ell$-th layer of $\bm{\phi}$, $0\leq\ell\leq L$. The Initial hidden embedding $\mathbf{h}^{0}$ is obtained by an MLP embedding layer that encodes the atom feature $\mathbf{h}]$. No embedding layer for the atom spatial coordinates. The construction of $\bm{\phi}_{\theta}$ is alternately updating the atom feature embeddings $\mathbf{h}$ and coordinates $\mathbf{x}$ as
@@
$\mathbf{x}_{i}^{\,\ell+1}$ $=\mathbf{x}_{i}^{\,\ell}+\sum_{\begin{subarray}{c}j\in\mathcal{V}(i)\\ j\neq i\end{subarray}}\bigl(\mathbf{x}_{i}^{\,\ell}-\mathbf{x}_{j}^{\,\ell}\bigr)\,f_{x}\bigl(d_{ij}^{\,\ell},\,\mathbf{h}_{i}^{\,\ell+1},\,\mathbf{h}_{j}^{\,\ell+1},\,\mathbf{e}_{ij};\,\theta_{x}\bigr)\,\mathbf{1}_{\mathrm{update}}.$ (10)
@@
A.2 E(3)-Equivariance Discussion
In molecular modeling, it is essential that the learned distribution over update nodes be invariant to translations, reflections, and rotations of the condition nodes. This E(3)-invariance reflects a fundamental inductive bias in molecular systems [15, 48]. Since SE(3) is a subgroup of E(3), any E(3)-equivariant model is also SE(3)-equivariant. While some prior works (e.g., [37, 7]) adopt the term SE(3), others (e.g., [15]) use E(3), which more accurately describes the symmetry of their networks. We follow the latter to avoid ambiguity.
The full Euclidean group in $\mathbb{R}^{3}$, denoted E(3), consists of rigid-body transformations of the form $T(x)=Rx+t$, where $R\in\mathbb{R}^{3\times 3}$ is a orthogonal matrix and $t\in\mathbb{R}^{3}$ is a translation vector.
If we pre-align the condition nodes by centering them at their center of mass (i.e., eliminating the translational degree of freedom), then the resulting likelihood becomes E(3)-invariant under the following condition: